If x and y are coordinates and t is time
x=f1(t)  -equ 1
y=f2(t)  -equ 2
Eliminating t we get y=f(x)
Pls explain how.
thanks
 

Grade:11

1 Answers

Himanshu Rathour
5 years ago
x=f1(t) 
This means that x is a function of t
y=f2(t)
This means that y is also a function of t
By eliminating t we can prove that y is a function of x
For eg take x=2t+3.... (i) 
y=3t..... (ii) 
From i & ii 
y=3/2(x-3)